OpenStudy (anonymous):

Maria earns $8 an hour babysitting. Last week Maria made $64. She needs to have at least $150 for vacation. Which inequality can be used to determine how many more hours Maria needs to babysit before she is ready for vacation?

OpenStudy (anonymous):

\[64x + 8 \ge 150\] \[64x + 8 \le 150\] \[8x + 64 \ge 150\] \[8x + 64 \le 150\]

OpenStudy (anonymous):

I think its C

OpenStudy (anonymous):

Any one help please?